OpenStudy (anonymous):

[Help me :3 Reward: Medal and Fan] The mean of 5 numbers is 18. If a number is excluded, the new mean changes to 16. Find the excluded number. Options : a) 24 b)28 c) 26 d) 29

OpenStudy (butterflydreamer):

5 numbers have an average of 18. So the total of the 5 numbers = 18*5 = 90 4 numbers have an average of 16 So the total of the 4 numbers = 16*4 = 64 You want to find the excluded number so subtract 90 - 64 to get your answer

OpenStudy (anonymous):

26? @butterflydreamer

OpenStudy (butterflydreamer):

yess

OpenStudy (anonymous):

You helped me atleast 3 times today! Thanks a lot! :)

OpenStudy (butterflydreamer):

haha no problemo :P!

OpenStudy (butterflydreamer):

please close the question if you're finished :)